OpenStudy (anonymous):

1lb of coal yields 1.6 x10000000 joules. Assuming a power plant is 30% efficient how much coal must be burned to light a 100 watt lightbulb for one day

OpenStudy (anonymous):

energy consume by 100 watt=100 joule/s light bulb in one day=100*24*3600 joule also let x lb coal we need so x lb coal give us energy x*(1.6 *10000000) as power plant is 30% efficient so [30% of x*(1.6 *10000000)] = 100*24*3600 joule i hope you can calculate x now.